OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of 2 and contains the point (4, 2)

OpenStudy (anonymous):

\[y-y_{1}=m(x-x _{1})\] y-2=2(x-4) y-2=2x-8 y=2x-6 Answer: y=2x-6